OpenStudy (anonymous):

absolute maximum and minimum of 6x^2 - 2x + 5 =

OpenStudy (anonymous):

6x^2 - 2x + 5, 0 =<x=<9

OpenStudy (anonymous):

vertex is \[-\frac{b}{2a}=\frac{1}{6}\] for the first coordinate, so minimum will be \[f(\frac{1}{6})\] max will be \[f(9)\]

OpenStudy (anonymous):

what are the x values also? or at which x values will you find the absolute min and max